40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 6 16:44:37 2011, steve, kiwamu, Update, SUS, BS oplev spectra BS.jpg
    Reply  Wed Jul 6 18:15:26 2011, kiwamu, Update, SUS, Re: BS oplev spectra 
    Reply  Tue Jul 12 17:30:24 2011, steve,, Update, SUS, BS oplev spectra BS_oplev.jpg
       Reply  Thu Jul 14 09:38:50 2011, steve,, Update, SUS, BS oplev spectra BS_oplev.jpg
          Reply  Thu Jul 14 15:27:08 2011, steve,, Update, SUS, SUS oplev spectras 7x
             Reply  Fri Jul 15 09:25:02 2011, rana, Update, SUS, SUS oplev spectras 
Message ID: 4948     Entry time: Wed Jul 6 18:15:26 2011     In reply to: 4947
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: Re: BS oplev spectra 

[Steve / Kiwamu]

Motivation:

 Since the oplevs were the ones we haven't carefully tested, so the oplevs need to be checked.

This checking is also a part of the suspension optimizations (see the minutes of the last 40m meeting).

 In this work Steve will check two things for all the oplevs :

    1. Noise level including the dark noise, electrical noise and ADC noise to just make sure that the noise are blow the signal levels below ~ 30Hz.

    2. The spectra of the signals to make sure there are no funny oscillations and unexpected structures

 

Measurement :

  To check the things listed above, we take two kinds of oplves' spectra :

     1. "dark noise" when the He-Ne beam is blocked.

     2. "signals" when the optics are damped by only OSEMs

 We did these checks on the BS oplev today (see the last entry).

All of them are fine, for example the dark noise (including electrical noise and ADC noise) are below the signal levels.

And no oscillation peak was found. Steve will go through all of the oplevs in this way.

Quote from #4947

Healthy BS oplev

ELOG V3.1.3-